# XRF V2: A Dataset for Action Summarization with Wi-Fi Signals, and IMUs in Phones, Watches, Earbuds, and Glasses

### v1

@@ 4.2. XRFMamba
Pyramidal Prediction Head. We use Mamba to transform the input into features with the same number of channels but different lengths, resulting in a feature pyramid composed of $(f_{1},f_{2},\dots,f_{K})$. Each feature dimension $f_{i}$ is processed through two separate convolutional paths to output action recognition results and action localization results, respectively. The total loss is the sum of the errors from these two prediction branches. The overall training loss of the network is given as follows:
(6) $L^{i}$ $=\alpha_{1}L_{Cls}^{i}+\alpha_{2}L_{Loc}^{i}$
$L$ $=\sum_{i=1}^{K}L^{i}$
Inference. In XRFMamba, each action has multiple localization priors, leading to several candidate prediction results. To convert multiple localization prediction results into a single prediction, we use Non-Maximum Suppression (NMS). The steps of NMS are as follows: (1) Sorting: sort all candidate segments by their scores, processing the segment with the highest score first. (2) Discard Overlapping Segments: after processing the current segment, compute its overlap with other segments. If the overlap exceeds a predefined threshold, that segment is discarded. (3) Repeat the Process: continue with the next highest-scoring segment and repeat step 3 until all segments have been processed. Fig. 7 shows an example. The segment corresponding to the highest classification score for action $i$ is selected first, and the other segments that have a high overlap with it are discarded. As a result, action $i$ will only retain one localization prediction. Then, among the remaining segments, the segment with the highest score for action $j$ is selected, and all the other segments with a high overlap with it are discarded. Therefore, action $j$ will also have only one prediction result. NMS enables XRFMamba to keep only the best result for each action.
4.3. Implementation Details
During the inference phase, XRFMamba processes 30-second temporal clips with a 5-second sliding stride, generating preliminary action proposals for each clip. To consolidate overlapping predictions, we implement the temporal non-maximum suppression (NMS) algorithm described above. The final temporal action localization (TAL) outputs comprise action classifications with precise start/end time boundaries.
5. Evaluation and Results
5.1. Metrics
(1) AP and mAP for Temporal Action Localization. We use the intersection over union (IoU) between the predicted action time and the ground truth time, referred to as tIoU, to evaluate the performance of Temporal Action Localization (TAL). A higher tIoU indicates better performance. tIoU can be computed as follows:
(7) $tIoU=\frac{time~{}of~{}overlap}{time~{}of~{}uniou}=\frac{\mathtt{max(0,\mathtt{min}(end_{pred},end_{gt}))}-\mathtt{max(0,\mathtt{min}(start_{pred},start_{gt}))}}{\mathtt{end_{pred}-start_{pred}+end_{gt}-start_{gt}}}$
where $\mathtt{start_{pred}}$ and $\mathtt{end_{pred}}$ are the predicted start and end times of an action; $\mathtt{start_{gt}}$ and $\mathtt{end_{gt}}$ are the ground truth start and end times.
(8) $AP@t=\frac{\sum_{i=1}^{N}\mathtt{I}(tIoU_{i}\geq t)}{N}$
where $\mathtt{Human}$ outputs 1 if the human auditor considers the LLM’s responses on ground truth and predicted action sequences share the same meaning. Specifically, for XRF V2, for the 171 test sequences, we prepared two prompts for each sequence, resulting in a total of 342 LLM responses to the ground truth and another 342 responses to the predicted results. All used prompts are presented in the Appendix. A.
We invited five auditors (3 male and 2 female), each with 1-3 years of experience using LLM models, to evaluate whether the meanings of the real and predicted responses were consistent for each of the 342 pairs. As shown in the formula, when the meanings are consistent, the numerator is incremented by 1, and the denominator represents the total number of evaluations, $N_{q}=342$. To reduce randomness, we perform action summarization using three different LLM models: ChatGPT-4o, Qwen2.5-Plus, and DeepSeek-V3, under their free plans. The final evaluation metric, mRMC, is obtained by averaging the results from different models and auditors. This provides a more robust and reliable measure for assessing the performance of action summarization.
5.2. Results
action mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg
Walk 99.65 99.33 99.33 99.33 95.20 91.55 87.82 72.56 52.50 45.95 84.32
mean 96.67 96.23 95.68 94.53 92.02 88.66 77.55 63.59 46.53 35.13 78.73
5.2.1. Temporal Action Localization Results
We evaluate several state-of-the-art approaches, including TriDet (Shi et al., 2023), ActionFormer (Zhang et al., 2022), and TemporalMaxer (Tang et al., 2023), which have been recently applied to IMU-based temporal action localization (Bock et al., 2024). We also compare with UWiFiAction (Wang et al., 2023), a WiFi-based action detection method, and WiFiTAD (Liu et al., 2025), a WiFi-based temporal action localization method. Additionally, we test ActionMamba (Chen et al., 2024), an advanced Mamba-based model for video-based action localization. All these methods are extended to support dual-input (IMU and WiFi) fusion, consistent with our approach. As shown in Table. 4, our method achieves the best performance, outperforming the recently accepted WiFiTAD (Liu et al., 2025) by a large margin of 5.49 points in mAP@avg, while using 35% fewer parameters.
We further present the mAP (mean Average Precision) for each action in Table. 5. From the table, it is clear that the mAP@avg for most actions remains relatively high, indicating that XRFMamba performs well in recognizing and localizing the majority of actions. However, some actions, such as Stretching and Opening Envelope, have relatively low mAP values compared to others. The Stretching action involves various body parts and features a wide range of motion with substantial variation in execution style. Different individuals may perform the stretching action in vastly different postures, which increases the complexity of recognizing the action accurately. In contrast, Opening an Envelope typically consists of subtle, short-duration hand movements, often involving just the fingers and palms. These localized and brief movements are more challenging to capture. Additionally, there is a sudden drop at 0.75 and 0.8 for Lying Still because it appeared only once in the test data, with its tIoUS being 0.7564.
The primary limitation of our method lies in the number of parameters that need to be learned, which is nearly 2.67 times that of ActionMamba (Chen et al., 2024). This is largely due to the use of TSSE (Temporal Signal Semantic Embedding) embedding module from WiFiTAD (Liu et al., 2025), which accounts for approximately 55.5M parameters. Therefore, a promising future research direction is to explore lighter and more efficient embedding modules that can maintain performance while significantly reducing parameters. Additionally, considering that the current state-of-the-art in video-based temporal action localization (TAL) method, such as ActionMamba (Chen et al., 2024), often relies on pre-trained models, such as InterVideo-6B (Wang et al., 2022a), as embedding modules, another potential research direction is to investigate the use of pre-trained time-series foundation models, such as TimesFM (Das et al., 2023), FoundTS (Li et al., 2024b), and Time-MoE (Shi et al., 2024), as the embedding module, thereby reducing the number of parameters that need to be learned during training.
5.2.2. Action Summarization Results
Table 6. XRFMamba achieves high response meaning consistency (RMC) in action summarization.
ChatGPT-4o Qwen2.5-Plus DeepSeek-V3 mean
RMC@a1 [278/342]0.813 [276/342]0.807 [278/342]0.813 0.811
RMC@a2 [276/342]0.807 [275/342]0.804 [276/342]0.807 0.806
RMC@a3 [279/342]0.816 [279/342]0.816 [276/342]0.807 0.813
RMC@a4 [278/342]0.813 [275/342]0.804 [277/342]0.810 0.809
RMC@a5 [262/342]0.766 [263/342]0.769 [266/342]0.778 0.771
mRMC 0.803 0.800 0.803 0.802
Figure 9. XRFMamba, when combined with LLMs, can function as an intelligent ambient sensing agent, such as personal assistant, health assistant, and home assistant, to provide services like task reminders, personalized health monitoring, climate control, and more.
We perform action summarization using three different LLM models: ChatGPT-4o, Qwen2.5-Plus, and DeepSeek-V3, all under their free plans. The consistency between predicted and real action summaries from five auditors is presented in Table 6. As shown in the table, auditors 1 to 4 consistently acknowledge that our LLMs’ predictions for TAL demonstrate high alignment with ground truth, achieving a response meaning consistency (RMC) of approximately 0.81. Auditor 5 exhibits slightly lower agreement with an RMC of 0.77. The mRMC across all auditors reaches 0.802, indicating XRFMamba achieves good TAL and action summarization results.
Notably, the auditor-specific RMC values show minimal correlation with LLM selection, indicating that XRFMamba, when integrated with free-plan LLMs, can effectively function as an intelligent ambient sensing agent. Practical implementations of such agents - including personal assistants, health monitoring systems, and smart home controllers (as illustrated in Fig. 9) - demonstrate capabilities ranging from task management and personalized health tracking to environmental regulation, thereby significantly enhancing smart home ecosystems through multifunctional intelligent services.
Fig. 10 presents prediction examples in Bedroom, Study room, and Dining room, with comparison methods including TemporalMaxer (Tang et al., 2023), ActionFormer (Zhang et al., 2022), TriDet (Shi et al., 2023), UWiFiAction (Wang et al., 2023), WiFiTAD (Liu et al., 2025), and ActionMamba (Chen et al., 2024) from Table. 4. It is visually clear that WiFiTAD, ActionMamba, and our XRFMamba perform better, with XRFMamba demonstrating superior classification accuracy and precise action boundary localization compared to state-of-the-art methods.
5.2.3. Leave-one-person-out Results
In real-world applications, XRFMamba should not be limited to specific individuals but should instead be adaptable to a wide range of users with varying characteristics, such as body movement patterns, walking styles, and device usage habits. To assess this adaptability to new users, we employ a leave-one-person-out cross-validation evaluation. In this evaluation, data from one volunteer is used as the test set, while data from all other users are used for training. As shown in Table 7, mAP for new users is 72.83, indicating that our method performs well and is capable of generalizing effectively to unseen users.
Table 7. Leave-one-person-out results. mAP for new users reaches 72.83, indicating that XRFMamba performs well and is capable of generalizing effectively to unseen users.
Test user mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg
user2 96.91 96.91 96.49 96.14 92.21 88.25 84.93 75.87 57.34 45.85 83.09
5.3. Ablation Study
5.3.1. How and Where to Conduct Fusion
XRFMamba takes both IMU and WiFi modalities from XRF V2 as input. We first test where and how modality fusion should be applied in the XRFMamba pipeline. We evaluate fusion after the projection, after embedding, and after the Mamba backbone. In terms of fusion strategies, we evaluate three methods: (1) simple weighted addition: applying fixed weight coefficients to the IMU (0.8) and Wi-Fi (0.2) features and directly adding the weighted features; (2) linear fusion: combining the IMU and Wi-Fi features through a learned linear transformation, followed by a gating mechanism using a sigmoid function to weight the contribution of each modality; and (3) gate fusion: applying separated convolutional layers to the IMU and Wi-Fi features to compute individual gating weights using a sigmoid function, which are then used to weight each modality and combine them through summation. For a detailed explanation of each fusion method, including the mathematical formulations and pseudocode, please refer to the Appendix. B.
The results of the nine different combinations of fusion positions and strategies are shown in Table 8. As illustrated in the table, fusion applied later in the pipeline involves more separate data streams and additional parameters. When comparing different fusion strategies, the gate and linear methods require more parameters than the weighted method, yet their performance is not as effective. In terms of fusion positions, combination 1, which applies weighted fusion after the projection module, achieves an average mAP of 78.74. In contrast, combination 3, which performs fusion after the Mamba backbone, achieves a higher mAP of 80.89 but requires nearly double the number of parameters. Therefore, we chose the weighted addition fusion method applied after the projection module as the final fusion approach for XRFMamba, as it offers the best mAP-parameter trade-off.
device mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg
1 95.88 95.32 93.52 90.90 85.02 76.06 60.50 40.57 17.40 2.93 65.81
2 94.07 93.03 91.17 87.33 79.99 67.60 51.39 27.82 10.97 1.75 60.51
3 80.99 79.94 77.65 74.81 71.11 64.19 53.78 37.90 19.02 2.86 56.22
4 91.91 90.38 85.54 79.48 68.81 54.26 35.51 18.48 7.98 1.16 53.35
5 85.76 83.82 80.99 76.34 70.48 56.36 43.89 29.39 13.03 2.76 54.28
6 89.06 88.14 86.48 83.64 75.66 66.33 50.27 30.88 9.77 1.42 58.16
7 65.40 60.51 54.67 47.75 34.70 24.68 13.17 5.52 1.67 0.32 30.84
8 95.58 95.07 93.77 88.44 85.46 77.62 67.22 49.08 26.28 4.91 68.34
9 95.19 94.60 92.87 90.86 86.38 78.39 65.48 46.95 18.51 4.58 67.38
10 91.03 89.70 87.84 82.18 78.23 71.02 57.76 37.97 18.55 4.59 61.89
11 92.40 91.80 90.77 87.64 84.49 76.19 65.69 49.87 28.32 7.66 67.48
12 94.86 94.02 93.50 91.69 89.33 82.84 71.61 52.37 22.11 3.06 69.54
13 95.29 94.60 93.57 92.19 89.82 83.28 68.06 44.47 18.34 2.76 68.24
14 94.36 94.15 92.62 90.66 87.59 80.33 68.98 51.75 31.19 10.64 70.23
15 93.05 92.76 91.91 90.15 86.76 81.01 69.77 54.25 37.55 17.17 71.44
16 96.22 95.16 94.43 92.24 89.17 79.57 68.14 47.16 22.77 4.58 68.94
17 96.36 95.87 95.26 93.23 90.65 87.08 75.10 59.85 38.74 11.36 74.35
18 96.82 96.51 96.23 95.31 92.84 84.06 73.08 54.41 26.99 3.95 72.02
19 96.51 95.82 95.44 94.23 92.00 86.98 80.26 62.57 41.40 29.00 77.42
20 97.48 97.33 96.77 92.02 89.02 84.29 76.69 60.04 41.98 18.64 75.43
21 96.67 96.23 95.68 94.53 92.02 88.66 77.92 63.96 46.53 35.13 78.74
5.3.2. Device Combination Study
XRF V2 dataset includes Wi-Fi signals and IMU data from seven types of devices: phones, watches, earbuds, and glasses, as shown in Fig.11. Users can combine devices according to their own needs. In Fig. 11, we report some of the most common device combinations, such as case 9 (a phone and a watch), case 14 (a phone, a watch, and an earbud), case 17 (a phone, a watch, glasses, and ambient Wi-Fi). It shows that the average mAP of using only a right-hand-worn smartwatch (case 1) is 65.81, and case 17 achieves an average mAP of 74.35, reaching a usable level.
6. Dataset and Code Availability, and Potential Usefulness
The data and code are available at https://github.com/aiotgroup/XRFV2. The repository includes a comprehensive $\mathtt{readme.md}$ file, which describes how to use the code to reproduce our results. The code features a well-structured modular design, making it easy to call and test various backbones and modality combinations. In this code repository, under the /model directory, you will find different method implementations, including XRFMamba and all comparison methods. You can easily use the model you need through make_model() found in /model/models.py. The code for XRFMamba is located in model/TAD_multi_fusion.py. Data loading methods are defined in the /dataset directory. To get started, you need to modify the paths in basic_config.json to match your system setup. Then, run script/train_run.py to train the model. Finally, you can evaluate the model by running script/test_run.py. For more details, please check out our repository.
We believe our dataset can significantly contribute to the community in at least the following areas.
  1. (1)
  2. (2)
  3. (3)
    XRF V2 is a dataset collected using real devices, and it provides high-quality data labels. It can be used to promote or evaluate research on IMU and WiFi data synthesis methods.
  4. (4)
  5. (5)
    XRF V2 supports multimodal learning between IMU, WiFi, and video, facilitating the development of multimodal foundational models. While the original videos cannot be shared due to privacy concerns with volunteers, we are to process them into features using InterVideo-6B (Wang et al., 2022a) and make these features publicly available.
  6. (6)
    We are to process the videos into 2D pose, 3D pose, and human mesh, which will support research in pose estimation or mesh reconstruction based on WiFi, IMU, and video data. Compared to previous datasets such as Person-in-WiFi 3D (Yan et al., 2024) and MM-Fi (Yang et al., 2024), where the person stands at a fixed position and repeats actions, the actions in XRF V2 are more representative of real-world scenarios, capturing more dynamic and varied human behavior.
  7. (7)
7. Conclusion
We present the XRF V2 dataset, which captures continuous action sequences from 16 volunteers across three typical household environments: the dining room, study room, and bedroom. Comprising a total of 825 action sequences, the dataset integrates data from various sensors, including IMUs from smartwatches, smartphones, earbuds, and glasses, Wi-Fi transceivers, and synchronized video from an Azure Kinect. The diverse and personalized action sequences, along with synchronized multimodal data, provide a valuable resource for advancing research in human action understanding and multimodal machine learning. We also introduce a novel human action understanding task: action summarization, for which we design a new evaluation metric, response meaning consistency (RMC). Action summarization takes action sequences and task-oriented prompts as input to large language models (LLMs), transforming the LLMs into intelligent agents. We demonstrate that with IMU or Wi-Fi input, LLMs can function as personal assistants, health assistants, and home assistants. Furthermore, we propose the XRFMamba neural network, which effectively performs temporal action summarization and action summarization, and conduct a detailed evaluation showing its superior performance compared to existing state-of-the-art methods.
Appendix A Evaluated Task-orientated Prompts
A.1. An example of responses from three LLMs to the same prompt.
@@ Appendix A Evaluated Task-orientated Prompts
A.2. All task-oriented questions used in action summarization evaluation.
@@
Appendix B Detailed Description of Three Fusion Methods

### v2

@@ Appendix C Decomposed Bidirectionally Mamba Block Process
Appendix D Statistical Significance Analysis
Table 11. Comparison of mAP performance for WiFiTAD and XRFMamba models
Model No. mAP@0.5 0.55 0.6 0.65 0.7 0.75 0.8 0.85 0.9 0.95 mAP@avg
WiFiTAD (Liu et al., 2025) 1 96.04 95.70 94.73 93.05 89.44 80.35 68.88 51.97 32.94 12.61 71.57
2 94.98 94.52 90.87 88.61 85.69 80.87 74.69 55.77 35.27 18.33 71.96
3 95.34 94.71 94.32 92.83 86.32 81.75 73.62 57.73 36.69 19.14 73.25
4 95.22 94.80 94.12 92.53 88.76 81.42 73.29 55.64 34.82 15.23 72.58
5 96.14 95.48 94.67 92.71 89.15 81.23 72.47 54.93 33.64 14.79 72.52
6 97.14 96.85 96.23 94.91 91.38 84.54 75.62 56.48 37.29 18.16 74.86
7 95.78 95.12 94.33 92.45 88.92 80.56 71.82 53.91 32.77 14.45 72.01
8 96.78 96.21 95.62 94.32 90.89 83.91 74.67 57.12 38.46 19.32 74.73
9 94.87 94.05 93.28 91.73 88.99 81.64 72.83 54.72 34.19 15.67 72.20
10 96.03 95.68 95.11 93.87 90.23 82.96 74.32 56.11 36.85 17.42 73.86
XRFMamba 1 97.62 96.90 95.90 95.40 93.18 89.78 82.02 65.87 47.06 36.13 79.99
2 97.53 97.21 96.89 96.47 93.54 90.82 80.36 64.11 48.30 38.41 80.36
3 96.83 96.44 96.19 95.54 93.27 88.88 82.20 68.53 51.55 40.23 80.97
4 97.50 97.15 96.79 95.83 92.69 89.70 78.40 63.70 46.13 36.22 79.41
5 96.53 96.04 95.83 95.05 92.13 87.05 80.67 66.51 51.27 38.30 79.94
6 95.56 94.59 93.82 93.61 91.08 86.64 80.77 62.79 48.55 38.50 78.59
7 96.97 96.54 96.03 93.04 89.75 84.69 77.59 64.75 49.79 38.31 78.75
8 96.49 96.23 93.01 92.00 89.55 84.88 78.25 64.22 46.95 35.18 77.68
9 97.26 96.87 96.70 95.73 91.95 87.18 81.14 67.23 47.90 37.00 79.89
10 97.04 96.55 96.35 95.61 92.82 88.86 79.81 65.78 50.71 38.24 80.18
Table 12. t-test Results Comparing XRFMamba and WiFiTAD at Different Thresholds
Threshold p-value Conclusion
mAP@0.5 $2.52\times 10^{-2}$ Significant
mAP@0.55 $4.00\times 10^{-2}$ Significant
mAP@0.6 $1.10\times 10^{-1}$ Not Significant
mAP@0.65 $4.16\times 10^{-2}$ Significant
mAP@0.7 $9.17\times 10^{-3}$ Significant
mAP@0.75 $1.31\times 10^{-4}$ Significant
mAP@0.8 $2.29\times 10^{-5}$ Significant
mAP@0.85 $4.42\times 10^{-7}$ Significant
mAP@0.9 $7.66\times 10^{-8}$ Significant
mAP@0.95 $3.57\times 10^{-10}$ Significant
mAP@avg $1.37\times 10^{-6}$ Significant
We conduct a detailed statistical significance analysis comparing our method (XRFMamba) with the state-of-the-art method (WiFiTAD). Each model is trained 10 times with random initialization (without using fixed seeds) to ensure the stability and reliability of the results, as summarized in Table 11. To determine whether the performance improvements of XRFMamba over WiFiTAD are statistically significant, we perform a paired sample t-test at each mAP@tIOU threshold (mAP@0.5, mAP@0.55, …, mAP@0.95, and mAP@avg). We compute the t-test using the following equation:
(10) $t=\frac{\bar{d}}{s_{d}/\sqrt{n}},$
where $d_{i}=x_{i}-y_{i}$ denotes the difference between paired results from XRFMamba ($x_{i}$) and WiFiTAD ($y_{i}$), $\bar{d}=\frac{1}{n}\sum_{i=1}^{n}d_{i}$ represents the mean of the paired differences, $s_{d}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(d_{i}-\bar{d})^{2}}$ is the standard deviation of these differences, and $n$ is the number of paired results (i.e., 10 repeated trials). We calculate the p-value based on the $t$-distribution with $n-1$ degrees of freedom, which quantifies the probability of observing such a difference (or a more extreme one) under the null hypothesis that the two methods perform equally. Generally, it is considered a statistically significant difference when $p<0.05$.
The results are presented in Table 12. For mAP@0.6, no significant difference is found between XRFMamba and WiFiTAD, with a p-value of 0.1103. This indicates that at this specific tIoU threshold, the performance of both methods is very similar, and there is insufficient evidence to claim a statistically significant difference. However, at most other thresholds (mAP@0.5, 0.55, 0.65, 0.7, 0.75, 0.8, 0.85, 0.9, and 0.95), the performance difference is statistically significant, with p-values consistently below 0.05. Notably, at higher tIoU thresholds, such as mAP@0.75, 0.8, 0.85, 0.9, and 0.95, the p-values are extremely low, demonstrating a strong and significant performance advantage of XRFMamba over WiFiTAD. Additionally, the mAP@avg result also shows a significant improvement, with a p-value of $1.37\times 10^{-6}$, further supporting the superior overall performance of XRFMamba. We hope these findings address your concerns and further strengthen the credibility of our results.
Appendix E Evaluated Task-orientated Prompts
E.1. An example of responses from three LLMs to the same prompt.
@@ Appendix E Evaluated Task-orientated Prompts
E.2. All task-oriented questions used in action summarization evaluation.
@@
Appendix F Detailed Description of Three Fusion Methods